# Diffusion Self-Weighted Guidance for Offline Reinforcement Learning

### v1

@@ 5.3 Ablation study for different weights
We then evaluated SWG’s performance under different weights formulations. We considered the Linex, Expectile and the exponential expectile weights presented in Sec. 4. This test was conducted on D4RL locomotion, using the same model and hyperparameter tuning to ensure a fair comparison, and the results are shown in Table 3. The expectile weight yielded the best performance, the exponential expectile showed moderate performance, and the Linex weight performed substantially worse. We conjecture that this is due to training instability introduced by the exponential weights in general. Most importantly, these results are consistent with the findings reported in IDQL idql .
@@ 5.4 How does SWG’s inference time scale with its size?
Since SWG requires computing gradients through the DM, we investigated its scalability properties via two experiments examining both depth and width scaling of its networks. For depth scaling, we used a simple multi-layer perceptron (MLP) with a hidden dimension of 2056 and measured the inference time across varying numbers of layers. For width scaling, we used a fixed 4-layer MLP with varying hidden dimension, measuring average inference time accordingly (See Appendix D.5).
As shown in Figures 3 and 3, SWG exhibited a linear scaling with respect to model depth and exponential scaling with respect to the width scaling (hidden dimension). Additionally, and as expected, sampling with guidance (i.e., computing gradients) translates to larger inference times. However, qualitatively, we can observe that both sampling methods scale at a similar rate. Note that scaling issues related to gradient computation can be easily avoided by designing the DM to output the weight component in its early layers. This allows gradients to be computed with respect to a fixed module, without requiring backpropagation through the entire model.
@@ 6 Conclusion
We have proposed a joint DM over actions and weights for optimal policy discovery in offline RL. Our approach addresses a shortcoming of current DM-based methods, which struggle to compute the guidance score network. In our proposal, the computation of the scores is straightforward, as they are readily available from the DM implementation. The proposed method, termed Self-Weighted Guidance (SWG), owes its name to having the weight guidance information embedded into the DM. Compared to previous methods, SWG conducts exact guidance using a streamlined setup which requires a single DM, and provides novel insights arising from modeling weights as random variables, thus demonstrating one more time the flexible modeling capabilities of DM’s.
Via toy examples and the standard D4RL benchmark, we have provided experimental validation of SWG’s ability to: i) effectively produce samples of the target distribution, and ii) perform comparably against state-of-the-art DMs on challenging environments such as Ant Maze, Franka Kitchen and Adroit Pen, while having a simpler training setup. Furthermore, we have provided an ablation study for different weight formulations within SWG, and also studied how SWG’s inference time scales with its network’s width and depth. We stress that the main advantage of SWG stems from its simplicity and reduced training setup compared with previous guidance methods.
Lastly, we claim that the proposed DM-based sampling methodology, which is guided by the same DM, has the potential for applications beyond the offline RL setting. Tasks involving weighted distributions such as controllable image and text generation diffusion_gans ; Li-2022-DiffusionLM can be addressed by the proposed SWG approach.
@@
Though SWG’s performed competitively on D4RL, it was outperformed by some recent diffusion methods released in the last two years, namely QGPO and D-DICE. We conjecture that this first version of SWG can be enhanced with tailored weight models and more sophisticated ODE solvers. Also, following our preliminary study on SWG’s scalability in Sec. 5.4 regarding gradient computation during inference, we claim that extracting the weight component from the first layers of the DM network will critically alleviate computational complexity.
@@
This work presents a promising methodology toward offline RL, where policies can be learned from a controlled offline dataset. However, as with most works on diffusion models, potential misuse includes the generation of harmful data, such as fake images, videos, or audio.
@@ Appendix A Proofs and derivations
We follow the derivation utilized in [31].
@@
[…]
@@
According to the diffusion models definition, we have:
[…]
@@
Recall that we want to sample from $p(\mathbf{z})$ in Eq. (18).
@@
To tackle the problem in Equation (10), let us assume we train a noise predicting diffusion model on the data $\mathbf{z}_{0}$, $\epsilon_{\theta}(\mathbf{z}_{k},k)$, $k\in[0,K]$, parametrized by $\theta$. Then, the optimal diffusion model $\epsilon^{*}$ is defined by the following optimization problem.
$\operatorname*{argmin}_{{\theta}}||{\epsilon}_{0}-{\epsilon}_{{\theta}}(\mathbf{z}_{k},k)||_{2}^{2}.$ (23)
Let us also recall the data prediction formula [8, 19] for diffusion models, which states that the optimal diffusion model satisfies:
@@
We can use this formula to compute the intractable score in Eq. (18). Since the extraction function $\phi_{w}(\cdot)$ is linear (see Def. 3.1 in Sec. 3), we can swap the expectation and $\phi_{w}$ to obtain
@@
Finally, plugging this term in Eq. (18)
@@
Appendix B Related work
In this section, we review prior work in key areas relevant to our approach, as a complement to Sec. 2, to better contextualize our contributions.
B.1 Offline RL
Offline RL methods must address common challenges such as overestimation errors for out-of-sample actions and distributional shift. Distributional shift is tackled by constraining the learned policy to remain close to the behavior policy, typically by adding a regularization term [25, 10, 35]. This regularization term can be utilized to learn the policy directly [10, 55, 25] or through a weighted regression procedure [41, 39, 35, 54].
Recent studies have proposed various methods for handling overestimation errors. For example, Conservative Q-Learning (CQL) [26] penalizes Q values of out-of-distribution actions to prevent overestimation. Implicit Q-Learning (IQL) [24] uses expectile regression to learn value functions, completely avoiding out-of-sample actions during training.
B.2 Diffusion models
Diffusion Models [48, 14, 50] are a recent powerful class of generative models that excel across numerous tasks, including, but not limited to, image synthesis [44, 7], text-to-image generation [43, 46] and video generation [16]. Controllable sampling has emerged as a necessary technique to ensure that generated samples possess desired features. Methods like Classifier Guidance [7] and Classifier-Free Guidance [15] aim to generate samples from the conditional distribution. Diffusion Models are limited by their slow sampling speed, as they require an autoregressive forward pass on a model. Existing attempts to improve sampling speed include: changing the Markovian formulation of Diffusion Models [49], distilling the model [47], and utilizing fast ODE solvers [33, 32].
B.3 Generative models in offline RL
Generative models have shown great promise in the field of Offline Reinforcement Learning (RL). Diffuser [17] performs planning with a diffusion model and improves results with a guidance network. Similarly, [1] proposes a method that utilizes conditional sampling to enhance results, employing Classifier-Free Guidance [15] to achieve this.
Diffusion policy methods have also been explored extensively. Diffusion-QL [53] adds Q-values as a regularization term during diffusion training to improve policy performance. Other notable methods include Implicit Diffusion Q-Learning (IDQL) [12] and SfBC [4] that model the behavior policy and simply utilize resampling from a batch of action candidates to improve performance. CEP [31], and D-DICE [34] are proposed as exact guidance methods, utilizing a diffusion model and an external guidance network to obtain samples from the target policy.
In addition to diffusion models, other generative models have been successfully applied to offline RL. Transformers, such as Decision Transformer (DT) [5] and Trajectory Transformer (TT) [18], have demonstrated impressive performance in various offline RL tasks. Variational Autoencoders (VAEs) [21] have also been employed to model the policy [10, 25].
Appendix C Relationship to previous guidance methods
In this section, we provide an analysis of previous guidance methods alongside our own, highlighting theoretical differences and potential advantages.
C.1 Comparison to previous guidance methods in offline RL
Let us consider the two existing exact-guidance DM methods in the offline RL literature: CEP [31] and IGL (from D-DICE [34]). Assuming a known weight function $w(\cdot)$, CEP learns the guidance network $g_{\theta}$ using the loss
where $k\sim\mathcal{U}(1,K)$, and for ease of notation $g^{(i)}_{\theta,k}=g_{\theta}(s,\mathbf{a}_{k}^{(i)},k)$, and $\mathbf{a}_{0,k}^{(1:M)}=\{\mathbf{a}_{j}^{(i)},\text{s.t.},j\in\{0,k\},i\in\{1,\ldots,M\}\}$. Recall that $\mathbf{a}_{0}$ is the data and $\mathbf{a}_{k}$ are the noisy copies. Since Eq. (28) requires $M>1$ action samples from each state in the dataset, its evaluation is likely to need out-of-sample actions, thus leading to overestimation errors [24, 34]. CEP [31] address this by pretraining a DM on the behavior policy and generating a fake action dataset containing $M>1$ samples for each state in the dataset.
On the other hand, IGL does not require querying out-of-sample actions using the following in-sample loss:
$\mathcal{L}_{IGL}(\theta)=\mathbb{E}_{k,\mathbf{a}_{0,k}}\Bigg{[}w(s,\mathbf{a}_{0})e^{-g_{\theta}(s,\mathbf{a}_{k},k)}+g_{\theta}(s,\mathbf{a}_{k},k)\Bigg{]},$ (29)
Though the objectives in Eqs. (28) and (29) produce unbiased estimates of the intractable term in Eq. (11), we argue that they either lack direct interpretability—as in the case of IGL—or are overly intricate. Also, they are challenging to work with due to the exponential terms $e^{g_{\theta}}$, which can lead to training instabilities and require additional considerations during model training. For instance, CEP addressed this by self-normalizing the exponential term, while IGL mitigated it by clipping the exponential.
$\mathcal{L}_{SWG}=\mathbb{E}_{k,\mathbf{z}_{0},\mathbf{z}_{k}}||{\epsilon_{0}}-{\epsilon}_{{\theta}}(\mathbf{z}_{k},k)||_{2}^{2},$ (30)
which is equivalent to maximizing a variational lower bound on the likelihood [19]. That is, a simple MSE over predicted noise terms free from exponentials prone to instabilities. Another relevant feature of SWG is that it inherently defines an in-sample loss, thus addressing the out-of-sample problem with the CEP objective.
Also, since SWG produces action and weights samples $\mathbf{z}_{0}=[\mathbf{a}_{0},w_{0}]$ simultaneously, it is possible to define a resampling mechanism as described in Sec 2.2 with the aim of further improving performance. However, our experimental results are obtained without utilizing resampling at all.
C.2 Comparison with conditional guidance
Notice that conducting conditional sampling with SWG is straightforward. Because for a label $c$ we can define the weight $w(\mathbf{a},c)\propto q(c|\mathbf{a})$, and recover the conditional sampling problem as $q(\mathbf{a}|c)\propto q(\mathbf{a})q(c|\mathbf{a})$. We claim that SWG and other exact weighted guidance methods [31, 34] generalize conditional sampling, as defining weights from label classes is straightforward—e.g., by training a classifier to estimate $q(c|\mathbf{a})$—whereas the reverse, deriving labels from weights, is not always feasible.
Appendix D Experimental details
D.1 2-D Toy examples
@@ Appendix D Experimental details
We followed the toy examples proposed in CEP [31], predefining a weight function $w(\mathbf{a})$ and a dataset distribution $q(\mathbf{a})$. We built the dataset $\mathcal{D}$ with 1M samples from $q(\mathbf{a})$ and extended it to $\mathcal{D}_{Z}=\{\mathbf{z}=(\mathbf{a},w(\mathbf{a}))\}_{i=1}^{N}$ using the predefined weight function.
@@
We then trained a diffusion model $\epsilon_{\theta}$ on the extended dataset $\mathcal{D}_{Z}$. For all toy examples, the diffusion model was trained for $K=100$ diffusion steps, using a Cosine noise schedule [36]. The architecture consisted of a 6-layer Multi-Layer-Perceptron (MLP) with hidden dimension 512 and GeLU activations. We trained over $200k$ gradient steps and batch size 1024 using the Adam optimizer [20] with learning rate $41e-$ 4.
@@
To avoid training a separate diffusion model for each $\beta$, we conditioned the diffusion model on $\beta$, $\epsilon_{\theta}(\mathbf{z}_{k},k,\beta)$, allowing us to train a single model per distribution $q(\mathbf{a})$. We trained over $\beta\in[0,20]$.
For the sampling process, we set the guidance scale $\rho$ to 1, as $\rho=1$ recovers the unbiased estimate of the target distribution score.
D.2 D4RL experiments
@@
For the D4RL Locomotion tasks, we averaged mean returns over 5 random seeds and 20 independent evaluation trajectories per seed. For Ant Maze tasks, we averaged over 50 evaluation trajectories and 5 random seeds. The results found in Table 2 correspond to the normalized score in each domain, following the suggestion of the dataset authors [9]. Our method results correspond to the evaluation at the last training step. The results for other methods are taken from their respective original articles, except for Franka Kitchen, QGPO and IDQL results were obtained from [28] and [37], respectively.
@@
Before training the models, we followed [9] and standardized locomotion task rewards in the dataset by dividing by the difference of returns of the best and worst trajectories in each dataset. For the Ant Maze dataset, we subtracted 1 from rewards.
@@
To perform the offline RL experiments, it was first necessary to estimate the weight function. We used the expectile weights presented in IDQL [12] for the main D4RL results, as we found that they produce the best performance compared to the tested weights.
@@
For the expectile weights, we needed to estimate the value functions $Q(s,\mathbf{a})$ and $V(s)$. We followed the same setup indicated in IQL [24] and IDQL [12], which consists of estimating Q-values $Q_{\psi}$ and the value function $V_{\phi}$, both with a 2-layer MLP with ReLU activations and 256 hidden units. We trained the networks with Adam optimizer [20] for 1M gradient steps, using learning rate $43e-$ 4 and batch size 256. We used double Q-learning [11] and soft updates [30] to stabilize the training. For the soft update, we used $0.005$ in all experiments to update the Q-network. We also set $\gamma=0.99$ for all tasks, and $\tau=0.7$ for the expectile loss, except for Ant Maze where we use $\tau=0.9$. The training procedure is described in Algorithm 3.
@@
Once the value functions are learned, we were able to calculate the weights as detailed in Sec. 4. Then, we extended our dataset to $\mathcal{D}_{Z}$ following the methodology described in Sec. 3. We trained a diffusion model $\epsilon_{\theta}$ on this extended dataset for $K=15$ diffusion steps in all experiments, using Adam optimizer [20] with learning rate $43e-$ 4 for all tasks, except for Adroit Pen where we use $53e-$ 5. We used a batch size of 1024, a variance-preserving noise schedule [50], and trained for 1M gradient steps in all tasks, except for Ant Maze, where we used 3M gradient steps.[17, 12]. The diffusion model consisted of a ResNet of 3 Residual blocks, with hidden dim 256 and Layer Normalization [2]. We used a dropout [51] of $0.1$ in all tasks. For the weight component output, we used the features from the middle layer of the ResNet and passed them through a simple MLP with a hidden dimension of 256 to obtain a more expressive representation of the weights. This architecture closely follows the one used in IDQL [12], with the exception of the additional MLP dedicated to the weight component. However, this MLP did not introduce a significant number of additional parameters. The full ResNet architecture contained approximately 1.8M parameters, ensuring a fair comparison in model capacity, as other guidance methods [31, 34] use U-Net architectures [45] with a comparable number of total parameters.
During inference, we tuned the guidance scale for each task. We swept over $\rho\in\{1,5,10,15,20,25,30\}$, such relatively large guidance scales were considered due to the expectile weights used, which were bounded in range. The guidance scale used for each experiment can be found in Tables 4, 5, 6 and 7. The results found in Table 2 correspond to the evaluation of the model at the last training step.
@@
Medium-Expert $15$ $5$ $5$
Medium $20$ $5$ $20$
@@
Play/Default $5$ $15$ $10$
Diverse $1$ $15$ $10$
@@
$1$ $25$
D.3 Ablation over guidance scale sensitivity
We provide an ablation study on using a fixed guidance scale across all tasks, without tuning $\rho$ for each task individually, to analyze how SWG and the expectile weights were affected by the guidance scale $\rho$. The choice of the guidance scale $\rho$ controls the strength of the guidance score, which mathematically translates into a sharper target distribution concentrated around high-value weights. This effect can be seen in the modified noise prediction $\hat{\epsilon}:={\epsilon}_{\theta}(\mathbf{z}_{k},k)-\rho\sigma_{k}\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}(\frac{\mathbf{z}_{k}-\sigma_{k}{\epsilon}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}))}={\epsilon}_{\theta}(\mathbf{z}_{k},k)-\sigma_{k}\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}((\frac{\mathbf{z}_{k}-\sigma_{k}{\epsilon}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}})^{\rho}))}$.
Overall, from Tables 8 and 9, we observe that the best performance was generally achieved with $\rho=10$ or $\rho=15$. As expected, using low guidance scales ($\rho<5$) resulted in suboptimal performance, as the guidance signal was not sufficiently strong. On the other hand, employing large guidance scales ($\rho>20$) led to a drop in performance, since an overly strong guidance signal tends to produce out-of-distribution samples.
Table 8: Guidance scale $\rho$ sensitivity in Locomotion tasks.
$\rho$ 1 5 10 15 20 25 30
Average 63.8 72.0 72.9 73.1 71.2 69.2 68.8
Table 9: Guidance scale $\rho$ sensitivity in Ant Maze tasks.
$\rho$ 1 5 10 15 20 25 30
Average 24.6 37.9 69.6 68.0 67.3 61.8 57.8
D.4 Ablation over the weights experiment
To ensure a fair comparison across weights, we used the same architecture with the same training hyperparameters. To find the best performance, we also tuned over the guidance scale.
Expectile weight: The expectile weight function was learnt with the same setup adopted in the main results, described in Sec. 4.
Exponential expectile weight: For the exponential expectile weight, we used the same setup as for IDQL expectile weights. And considered a temperature $\beta=3$ to exponentiate the advantage function $e^{\beta{A(s,\mathbf{a})}}$. To ensure stability during training, we clipped the exponential weights to $(0,80]$.
Linex weight: We followed IDQL [12], and utilized $\alpha=1$ to conduct the experiments.
D.5 Scaling of inference time in SWG
Each data point in Figures 3 and 3 represents the average inference time with our method for $K=15$ diffusion steps, obtained over 10,000 runs of the algorithm, along with the corresponding error region. For completeness, we report the total number of parameters in each network used in our experiments. The table below shows the parameter counts corresponding to different width and depth configurations:

### v2

@@ 6.3 Ablation study for different weights
We then evaluated SWG’s performance under different weight formulations in Sec. 5. This test was conducted on D4RL locomotion, using the same model architecture and hyperparameter tuning to ensure a fair comparison. Furthermore, we also tuned the guidance scale in each case to find the best possible performance. The weight formulations were implemented as follows:
  • Expectile weight: Learnt with the same setup adopted in the main results, described in Sec. 5.
  • Exponential expectile weight: We used the same setup as for IDQL expectile weights, and considered a temperature $\beta=3$ to exponentiate the advantage function $e^{\beta{A(s,\mathbf{a})}}$. To ensure stability during training, we clipped the exponential weights to $(0,80]$.
  • Linex weight: We followed IDQL (idql), and used $\alpha=1$.
The results, shown in Table 3, reveal that expectile yielded the best performance, with the exponential expectile showing moderate performance, and lastly Linex performing substantially worse. We conjecture that this is due to the training instability introduced by the exponential weights, in general. Most importantly, these results are consistent with the findings reported in IDQL (idql).
@@
6.4 Ablation over guidance scale sensitivity
We also studied the effect of using a fixed guidance scale across all tasks, without tuning $\rho$ for each task individually. Recall that the choice of the guidance scale $\rho$ controls the strength of the guidance score, which translates into a sharper target distribution concentrated around high-value weights. This effect can be seen in the modified noise prediction $\hat{\epsilon}:={\epsilon}_{\theta}(\mathbf{z}_{k},k)-\rho\sigma_{k}\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}(\frac{\mathbf{z}_{k}-\sigma_{k}{\epsilon}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}))}={\epsilon}_{\theta}(\mathbf{z}_{k},k)-\sigma_{k}\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}((\frac{\mathbf{z}_{k}-\sigma_{k}{\epsilon}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}})^{\rho}))}$. We conduct these experiments using expectile weights, which, as shown in Sec. 6.3, yield superior performance.
Overall, from Tables 4 and 5, we observe that the best performance was generally achieved with $\rho=10$ or $\rho=15$. As expected, using low guidance scales ($\rho<5$) resulted in suboptimal performance, as the guidance signal was not sufficiently strong. On the other hand, employing large guidance scales ($\rho>20$) led to poorer performance, since an overly strong guidance signal tends to produce out-of-distribution samples.
Table 4: Guidance scale $\rho$ sensitivity in Locomotion tasks.
$\rho$ 1 5 10 15 20 25 30
Average 63.8 72.0 72.9 73.1 71.2 69.2 68.8
Table 5: Guidance scale $\rho$ sensitivity in Ant Maze tasks.
$\rho$ 1 5 10 15 20 25 30
Average 24.6 37.9 69.6 68.0 67.3 61.8 57.8
@@ 6.5 How does SWG’s inference time scale with its size?
Since SWG requires computing gradients through DM, we investigated its scalability properties through two experiments examining both depth and width scaling of its networks. For depth scaling, we used a simple multilayer perceptron (MLP) with a hidden dimension of 2056 and measured the inference time across varying numbers of layers. For width scaling, we used a fixed 4-layer MLP with varying hidden dimension, measuring the average inference time accordingly (see the Appendix B.3).
As shown in Figures 5 and 5, SWG exhibited a linear scaling with respect to model depth and exponential scaling with respect to the width scaling (hidden dimension). Additionally, and as expected, sampling with guidance (i.e., computing gradients) translates into larger inference times. However, qualitatively, we can observe that both sampling methods scale at a similar rate. Note that scaling issues related to gradient computation can be easily avoided by designing the DM to output the weight component in its early layers. This allows gradients to be computed with respect to a fixed module, without requiring backpropagation through the entire model. Consequently, the additional computational overhead of SWG remains constant, even as the sub-component of the model responsible for action generation increases in scale.
The training time remains essentially the same as that of the related methods (idql; diff_dice) when using a similar architecture. The SWG training time measurements are reported in the Appendix B.4.
@@ 7 Conclusion
We have proposed a joint DM over actions and weights for optimal policy discovery in offline RL. Unlike current DM-based methods, our approach avoids the computation of a guidance score network. In our proposal, the computation of the scores is straightforward, as they are readily available from the DM implementation. The proposed method, termed Self-Weighted Guidance (SWG), owes its name to having the weight guidance information embedded into the DM. Compared to previous methods, SWG conducts exact guidance using a streamlined setup which requires a single DM, and provides novel insights arising from modeling weights as random variables, thus demonstrating one more time the flexible modeling capabilities of DM’s.
Via toy examples and the standard D4RL benchmark, we have provided experimental validation of SWG’s ability to: i) effectively produce samples of the target distribution, and ii) perform on par with —or better than— the state-of-the-art DMs on challenging environments such as Ant Maze, Franka Kitchen and Adroit Pen, while having a simpler training setup. We have also validated the ability of SWG to leverage resampling for improved performance. However, we observe that, in general, SWG performs sub optimally on mixed-action datasets (i.e., Mixed, Medium-Replay, Diverse), which is likely due to either the choice of weight function and/or the architecture of the DM, both of which may limit how effectively the joint action–weight distribution is learned. Furthermore, we have provided an ablation study for different weight formulations within SWG, and also studied how SWG’s inference time scales with its network’s width and depth. We stress that the main advantage of SWG comes from its simplicity and reduced training setup compared to previous guidance methods.
Lastly, we claim that the proposed DM-based sampling methodology, which is guided by the same DM, has the potential for applications beyond the offline RL setting. Tasks involving weighted distributions, such as controllable image and text generation (diffusion_gans; Li-2022-DiffusionLM), can be addressed by the proposed SWG approach.
@@
We have provided a proof of concept for SWG, which performed among the top 2 methods across D4RL’s datasets except in Locomotion, where SWG was outperformed by QGPO, D-QL and D-DICE. To address intrinsic limitations, we note that jointly learning actions and weights reduces flexibility (changing the guidance requires retraining the entire diffusion model) and makes training more challenging compared to learning actions alone, potentially needing higher-capacity networks. Both issues could be mitigated by dedicating a portion of the network exclusively to learning weights. Future work could further improve SWG with tailored weight models and more sophisticated ODE solvers.
@@
This work presents a promising methodology for offline RL, where policies can be learned from a controlled offline dataset. However, as with most diffusion models, potential misuse includes the generation of harmful data, such as fake images, videos, or audio.
Acknowledgements
JRdS acknowledges financial support from ANID-Chile grants Basal CIA250010 and Fondecyt-Regular 1251823.
@@ Appendix A Proofs and derivations
We follow the derivation utilized in (cep).
@@
$A=\int q_{0}(\mathbf{z}_{0})\phi_{w}(\mathbf{z}_{0})d\mathbf{z}_{0}=\mathbb{E}_{q_{0}(\mathbf{z}_{0})}\left[\phi_{w}(\mathbf{z}_{0})\right].$
@@
According to the definition of diffusion models, we have:
$p_{k}(\mathbf{z}_{k})=\int p_{k0}(\mathbf{z}_{k}|\mathbf{z}_{0})p_{0}(\mathbf{z}_{0})d\mathbf{z}_{0}=\int p_{k0}(\mathbf{z}_{k}|\mathbf{z}_{0})q_{0}(\mathbf{z}_{0})\frac{\phi_{w}(\mathbf{z}_{0})}{A}d\mathbf{z}_{0}=\frac{1}{A}\int q_{k0}(\mathbf{z}_{k}|\mathbf{z}_{0})q_{0}(\mathbf{z}_{0})\phi_{w}(\mathbf{z}_{0})d\mathbf{z}_{0}$
$q_{k}(\mathbf{z}_{k})\frac{1}{A}\int q_{0}(\mathbf{z}_{0}|\mathbf{z}_{k})\phi_{w}(\mathbf{z}_{0})d\mathbf{z}_{0}=q_{k}(\mathbf{z}_{k})\frac{\mathbb{E}_{q_{0}(\mathbf{z}_{0}|\mathbf{z}_{k})}\left[\phi_{w}(\mathbf{z}_{0})\right]}{A},$
@@
Recall that we want to sample from $p(\mathbf{z})$ in equation 10.
@@
To address the problem in Equation equation 4, let us assume that we train a noise-predicting diffusion model on the data $\mathbf{z}_{0}$, $\epsilon_{\theta}(\mathbf{z}_{k},k)$, $k\in[0,K]$, parametrized by $\theta$. Then, the optimal diffusion model $\epsilon^{*}$ is defined by the following optimization problem.
[…]
Let us also recall the data prediction formula (efron2011tweedie; variational_diffusion_models) for diffusion models, which states that the optimal diffusion model satisfies:
@@
We can use this formula to compute the intractable score in equation 10. Since the extraction function $\phi_{w}(\cdot)$ is linear (see Def. 3.1 in Sec. 3), we can swap the expectation and $\phi_{w}$ to obtain
@@
Finally, plugging this term in equation 10
@@
Appendix B Experimental details
B.1 2-D Toy examples
@@ Appendix B Experimental details
We followed the toy examples proposed in CEP (cep), predefining a weight function $w(\mathbf{a})$ and a dataset distribution $q(\mathbf{a})$. We built the dataset $\mathcal{D}$ with 1M samples from $q(\mathbf{a})$ and extended it to $\mathcal{D}_{Z}=\{\mathbf{z}=(\mathbf{a},w(\mathbf{a}))\}_{i=1}^{N}$ using the predefined weight function.
@@
We then trained a diffusion model $\epsilon_{\theta}$ on the extended dataset $\mathcal{D}_{Z}$. For all toy examples, the diffusion model was trained for $K=100$ diffusion steps, using a Cosine noise schedule (nichol2021improveddenoisingdiffusionprobabilistic). The architecture consisted of a 6-layer Multi-Layer-Perceptron (MLP) with hidden dimension 512 and GeLU activations. We trained over $200k$ gradient steps and batch size 1024 using the Adam optimizer (kingma2014adam) with learning rate $10^{-4}$.
@@
To avoid training a separate diffusion model for each $\beta$, we conditioned the diffusion model on $\beta$, $\epsilon_{\theta}(\mathbf{z}_{k},k,\beta)$, allowing us to train a single model per distribution $q(\mathbf{a})$. We trained over $\beta\in[0,20]$. For the sampling process, we set the guidance scale $\rho$ to 1, as $\rho=1$ recovers the unbiased estimate of the target distribution score.
Figure 6 presents four additional experiments on the swiss roll, rings, moons, and checkerboard distributions.
Figure 6: Additional examples for the experiment in Sec. 6.1. Target distributions from top-left, clockwise: swiss roll, rings, moons, and checkerboard distributions
B.2 D4RL experiments
@@
For the D4RL Locomotion tasks, we averaged mean returns over 5 random seeds and 20 independent evaluation trajectories per seed. For Ant Maze tasks, we averaged over 50 evaluation trajectories and five random seeds. The results found in Table 2 correspond to the normalized score in each domain, following the suggestion of the authors of the dataset (fu2020d4rl). The results of our method correspond to the evaluation in the last training step. The results for other methods are taken from their respective original articles, except for Franka Kitchen, QGPO and IDQL results were obtained from (li2024efficient_ld) and (fql_park2025), respectively.
@@
Before training the models, we followed the suggestions found in (fu2020d4rl) and standardized locomotion task rewards in the datasets by dividing them by the difference of returns of the best and worst trajectories in each dataset. For the Ant Maze dataset, we subtracted 1 from the rewards.
@@
To perform the offline RL experiments, it was first necessary to estimate the weight function. We used the expectile weights presented in IDQL (idql) for the main D4RL results, as we found that they produce the best performance compared to the weights tested.
@@
For the expectile weights, we needed to estimate the value functions $Q(s,\mathbf{a})$ and $V(s)$. We followed the same setup indicated in IQL (iql) and IDQL (idql), which consists of estimating Q-values $Q_{\psi}$ and the value function $V_{\phi}$, both with a 2-layer MLP with ReLU activations and 256 hidden units. We trained the networks with Adam optimizer (kingma2014adam) for 1M gradient steps, using learning rate 3e43e-4 and batch size 256. We used double Q-learning (fujimoto_dqn) and soft updates (soft_updates) to stabilize the training. For the soft update, we used $0.005$ in all experiments to update the Q-network. We also set $\gamma=0.99$ for all tasks, and $\tau=0.7$ for the expectile loss, except for Ant Maze where we use $\tau=0.9$. The training procedure is described in Algorithm 3.
@@
Once the value functions are learned, we were able to calculate the weights as detailed in Sec. 5. Then, we extended our dataset to $\mathcal{D}_{Z}$ following the methodology described in Sec. 3. We trained a diffusion model $\epsilon_{\theta}$ on this extended dataset for $K=15$ diffusion steps in all experiments, using Adam optimizer (kingma2014adam) with learning rate 3e43e-4 for all tasks, except for Adroit Pen where we use 3e53e-5. We used a batch size of 1024, a variance-preserving noise schedule (songscore), and trained for 1M gradient steps in all tasks, except for Ant Maze, where we used 3M gradient steps.(janner2022diffuser; idql). The diffusion model consisted of a ResNet of 3 Residual blocks, with hidden dim 256 and Layer Normalization (layernorm). We used a dropout (dropout) of $0.1$ in all tasks. For the weight component output, we used the features from the middle layer of the ResNet and passed them through a simple MLP with a hidden dimension of 256 to obtain a more expressive representation of the weights. This architecture closely follows the one used in IDQL (idql), with the exception of the additional MLP dedicated to the weight component. However, this MLP did not introduce a significant number of additional parameters. The complete ResNet architecture contained approximately 1.8M parameters, ensuring a fair comparison in model capacity, as other guidance methods (cep; diff_dice) use U-Net architectures (unet) with a comparable number of total parameters.
During inference, we tuned the guidance scale for each task. We swept over $\rho\in\{1,5,10,15,20,25,30\}$, such relatively large guidance scales were considered due to the expectile weights used, which were bounded in range. The guidance scale used for each experiment to obtain SWG results can be found in Tables 6, 7, 8 and 9. The results found in Table 2 correspond to the evaluation of the model at the last training step. For SWG-R (with the added resampling step), we conducted an additional hyperparameter search over the batch size, sweeping $M\in\{1,2,4,8,16\}$.
@@
Medium-Expert $15$ 55 55
Medium 2020 55 2020
@@
Play/Default 55 $15$ $10$
Diverse 11 $15$ $10$
@@
11 $25$
B.3 Scaling of inference time in SWG
Each data point in Figures 5 and 5 represents the average inference time with our method for $K=15$ diffusion steps, obtained over 10,000 runs of the algorithm, along with the corresponding error region. For completeness, we report the total number of parameters in each network used in our experiments. The table below shows the parameter counts corresponding to different width and depth configurations: